# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v3

@@ 3.4 Pairwise-Canceling Masks
The key innovation of our protocol is the use of pairwise-canceling masks. For any signing set $S$, each party $i\in S$ computes a mask $\mathbf{m}_{i}\in R_{q}^{\ell}$ such that $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$.
@@
where $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{d}$ is instantiated with SHAKE-256, and $d$ is the appropriate dimension ($\ell$ for response masks, $k$ for commitment masks).
We use domain separators $\mathsf{dom}=\mathsf{nonce}\|\texttt{"resp"}$ for response masks $\mathbf{m}_{i}\in R_{q}^{\ell}$ and $\mathsf{dom}=\mathsf{nonce}\|\texttt{"comm"}$ for commitment masks $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$.
@@ 3.5 Signing Protocol
The signing protocol is a three-round interactive protocol among parties in the signing set $S$ (where $|S|\geq T+1$ to ensure at least two honest parties; see Remark 4.1). We assume a broadcast channel or a designated combiner (TEE/HSM in Profile P1).
@@
4:Round 1: Nonce Commitment
5:for each party $i\in S$ in parallel do
6:  Sample $\mathbf{y}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$
7:  Compute $\mathbf{w}_{i}=\mathbf{A}\mathbf{y}_{i}$
8:  Sample $r_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
9:  Compute $\mathsf{Com}_{i}=H(\texttt{"com"}\|\mathbf{y}_{i}\|\mathbf{w}_{i}\|r_{i})$
10:  Broadcast $\mathsf{Com}_{i}$
11:  Store $(\mathbf{y}_{i},\mathbf{w}_{i},r_{i})$ as local state
12:end for
13:
14:Round 2: Masked Reveal and Challenge
15:Pre-round: Derive preliminary nonce $\mathsf{nonce}_{0}=H(\texttt{"nonce0"}\|\mathsf{Com}_{1}\|\cdots\|\mathsf{Com}_{|S|})$
16:for each party $i\in S$ in parallel do
17:  Compute commitment mask $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$ using $\mathsf{nonce}_{0}\|\texttt{"comm"}$
18:  Broadcast $(\mathbf{W}_{i}=\mathbf{w}_{i}+\mathbf{m}_{i}^{(w)},r_{i})$ $\triangleright$ Masked commitment
19:end for
20:All parties / Combiner:
21: Aggregate $\mathbf{W}=\sum_{j\in S}\mathbf{W}_{j}=\sum_{j}\mathbf{w}_{j}+\sum_{j}\mathbf{m}_{j}^{(w)}=\mathbf{w}$ $\triangleright$ Masks cancel
22: Compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{W},2\gamma_{2})$
23: Compute $\tilde{c}=H_{\mathsf{chal}}(\mu\|\mathbf{w}_{1})$ and expand to challenge $c\in R_{q}$ $\triangleright$ FIPS 204 spec
24: Derive $\mathsf{nonce}=H(\texttt{"nonce"}\|c\|\mu\|\mathsf{sort}(S))$
25:
26:Round 3: Masked Response
27:for each party $i\in S$ in parallel do
28:  Compute Lagrange coefficient $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\mod q$
29:  Compute masked secret contribution: $\boldsymbol{\nu}_{i}=\lambda_{i}\cdot c\cdot\mathbf{s}_{1,i}$
30:  Compute mask $\mathbf{m}_{i}$ as in Definition 3.1
31:  Compute masked response: $\mathbf{U}_{i}=\mathbf{y}_{i}+\boldsymbol{\nu}_{i}+\mathbf{m}_{i}$
32:  Compute $\mathbf{V}_{i}=\lambda_{i}\cdot c\cdot\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$ $\triangleright$ For r0-check
33:  Send $(\mathbf{U}_{i},\mathbf{V}_{i})$ to combiner
34:end for
35:
36:Aggregation and r0-Check (see Subsection 3.9)
37:Combiner computes $\mathbf{z}=\sum_{i\in S}\mathbf{U}_{i}\mod q$
38:if $\left\|\mathbf{z}\right\|_{\infty}\geq\gamma_{1}-\beta$ then
39:  return $\bot$ (z-bound abort, retry from Round 1)
40:end if
41:r0-check: Verify $\left\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\right\|_{\infty}<\gamma_{2}-\beta$
42:if r0-check fails then
43:  return $\bot$ (r0 abort, retry from Round 1)
@@
45:Compute $\mathbf{r}=\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d}$ $\triangleright$ $\mathbf{w}-c\mathbf{s}_{2}+c\mathbf{t}_{0}$
46:Compute $c\mathbf{t}_{0}$ using $c\mathbf{s}_{2}$ from r0-check $\triangleright$ See Subsection 3.9
47:Compute hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{r},2\gamma_{2})$
48:if $\mathsf{weight}(\mathbf{h})>\omega$ then
49:  return $\bot$ (hint weight abort)
50:end if
51:return signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
@@
The Round 1 commitments $\mathsf{Com}_{i}$ bind parties to their nonces before seeing others’ values, preventing adaptive attacks. However, we do not publicly verify these commitments during normal signing—the masked values $\mathbf{W}_{i}$ and $\mathbf{U}_{i}$ cannot be verified without revealing private mask seeds. Instead, we follow an optimistic model: if signing fails (excessive aborts, invalid signatures), the TEE coordinator triggers a blame protocol where parties reveal their commitments for verification. This “commit always, open only on blame” approach minimizes latency on the normal path while enabling accountability. See Section 6.2 for details.
@@
We use distinct domain tags ("com", "nonce0", "nonce", "resp", "comm") for all hash function uses to ensure random oracle independence.
@@
When $\gamma_{1}$ is not divisible by $|S|$, rounding slightly reduces the nonce range. This has negligible impact: for ML-DSA-65 with $|S|=17$, the range loss is $2\times 10^{-5}$ per coordinate and does not affect the z-bound check since $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor>\gamma_{1}-\beta$. See Appendix G for details.
@@ 3.6 Correctness
Proof 3.8.
The aggregated response is $\mathbf{z}=\sum_{i\in S}\mathbf{U}_{i}=\sum_{i}\mathbf{y}_{i}+c\cdot\sum_{i}\lambda_{i}\mathbf{s}_{1,i}+\sum_{i}\mathbf{m}_{i}=\mathbf{y}+c\mathbf{s}_{1}$, using nonce summation, Lagrange reconstruction, and mask cancellation (Lemma 3.2). Since $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$, verification proceeds as in single-signer ML-DSA.
@@
(1) While individual $\mathbf{U}_{i}$ may have large coefficients, the sum $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ is short (<q/2<q/2), so mod $q$ reduction is correct. (2) Party $i$ must compute $\mathbf{U}_{i}=\mathbf{y}_{i}+\lambda_{i}\cdot c\cdot\mathbf{s}_{1,i}+\mathbf{m}_{i}$, not $\lambda_{i}\cdot(\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i})+\mathbf{m}_{i}$ (the latter incorrectly weights nonces). (3) Signatures are FIPS 204-compatible: same format, size (3309 bytes for ML-DSA-65), and pass unmodified verifiers. Signer-side differences (Irwin-Hall nonces, interactive protocol) do not affect security or verification.
@@ 3.7 Nonce Distribution and Security
The aggregated nonce $\mathbf{y}=\sum_{i\in S}\mathbf{y}_{i}$ follows an Irwin-Hall distribution concentrated around zero. This preserves EUF-CMA security: using smooth Rényi divergence bounds from [Raccoon2024], we verify $R^{\epsilon}_{2}\leq 1.003$ for $|S|\leq 17$ and $R^{\epsilon}_{2}\leq 1.05$ for $|S|\leq 33$ signers (Appendix D), yielding negligible security loss ($2^{-40}$ even for $2^{64}$ queries).
@@ 4.1 Security Model
Remark 4.1 (Why $|S|\geq T+1$: Two-Honest Requirement).
Our protocol requires signing sets of size at least $T+1$ rather than $T$. This ensures that even when the adversary corrupts the maximum |C|=T1|C|=T-1 parties, every signing set contains at least two honest parties—necessary for mask privacy (see Lemma 4.6).
@@ 4.2 Unforgeability
$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}$
where $q_{H},q_{s}$ are random oracle and signing queries. The security loss matches single-signer ML-DSA [AFLT12, DKL18].
@@
Simulate signing without knowing $\mathbf{s}_{1}$: sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1}+\beta+1,\ldots,\gamma_{1}-\beta-1\}^{n\ell}$, compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$, and program $H(\mu\|\mathbf{w}_{1}):=c$. The masked values $\mathbf{W}_{j},\mathbf{U}_{j}$ are simulated as uniform by Lemma 4.6. By PRF security: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3}]|\leq\binom{|S|}{2}\cdot\epsilon_{\mathsf{PRF}}$.
@@
This yields a short vector for the augmented M-SIS instance $\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]$ with $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),\tau)<2\gamma_{1}$. Thus: $\Pr[\mathsf{Win}_{4}]\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}$.

### v4

@@ 3.5 Pairwise-Canceling Masks
Pairwise-canceling masks are used to hide the commitment values $\mathbf{W}_{i}$ (for challenge derivation) and the r0-check shares $\mathbf{V}_{i}$ (which contain $\lambda_{i}\cdot c\cdot\mathbf{s}_{2,i}$). For any signing set $S$, each party $i\in S$ computes a mask $\mathbf{m}_{i}$ such that $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$.
@@
where $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{d}$ is instantiated with SHAKE-256, and $d$ is the appropriate dimension ($k$ for commitment masks, $k$ for r0-check masks).
We use domain separators $\mathsf{dom}=\mathsf{nonce}\|\texttt{"comm"}$ for commitment masks $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$ and $\mathsf{dom}=\mathsf{nonce}\|\texttt{"s2"}$ for r0-check masks $\mathbf{m}_{i}^{(s2)}\in R_{q}^{k}$.
@@ 3.6 Signing Protocol
The signing protocol is a three-round interactive protocol among parties in the signing set $S$ (where $|S|\geq T$), preceded by an offline nonce DKG preprocessing step (Algorithm 2). We assume a broadcast channel or a designated combiner (TEE/HSM in Profile P1).
@@
4:Preprocessing: Nonce DKG (offline)
5:Execute $\mathsf{NonceDKG}(S)$ (Algorithm 2) to obtain nonce shares $\{\mathbf{y}_{i}\}_{i\in S}$
6:
7:Round 1: Nonce Commitment
8:for each party $i\in S$ in parallel do
9:  Compute $\mathbf{w}_{i}=\mathbf{A}\mathbf{y}_{i}$ $\triangleright$ Using DKG nonce share
10:  Sample $r_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
11:  Compute $\mathsf{Com}_{i}=H(\texttt{"com"}\|\mathbf{y}_{i}\|\mathbf{w}_{i}\|r_{i})$
12:  Broadcast $\mathsf{Com}_{i}$
13:  Store $(\mathbf{y}_{i},\mathbf{w}_{i},r_{i})$ as local state
14:end for
15:
16:Round 2: Masked Reveal and Challenge
17:Pre-round: Derive preliminary nonce $\mathsf{nonce}_{0}=H(\texttt{"nonce0"}\|\mathsf{Com}_{i_{1}}\|\cdots\|\mathsf{Com}_{i_{|S|}})$ where $(i_{1},\ldots,i_{|S|})=\mathsf{sort}(S)$
18:for each party $i\in S$ in parallel do
19:  Compute Lagrange coefficient $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\bmod q$
20:  Compute commitment mask $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$ using $\mathsf{nonce}_{0}\|\texttt{"comm"}$
21:  Send $(\mathbf{W}_{i}=\lambda_{i}\cdot\mathbf{w}_{i}+\mathbf{m}_{i}^{(w)},r_{i})$ to coordinator $\triangleright$ NOT broadcast; see Remark 4.1
22:end for
23:Coordinator:
24: Aggregate $\mathbf{W}=\sum_{j\in S}\mathbf{W}_{j}=\sum_{j}\lambda_{j}\mathbf{w}_{j}+\sum_{j}\mathbf{m}_{j}^{(w)}=\mathbf{A}\mathbf{y}=\mathbf{w}$ $\triangleright$ Masks cancel
25: Compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{W},2\gamma_{2})$
26: Compute $\tilde{c}=H_{\mathsf{chal}}(\mu\|\mathbf{w}_{1})$ and expand to challenge $c\in R_{q}$ $\triangleright$ FIPS 204 spec
27:if $c$ is not invertible in $R_{q}$ then
28:  return $\bot$ (challenge invertibility abort, retry from Round 1) $\triangleright$ Negligible probability
29:end if
30: Derive $\mathsf{nonce}=H(\texttt{"nonce"}\|\tilde{c}\|\mu\|\mathsf{sort}(S))$
31: Broadcast $(\mathbf{w}_{1},\tilde{c},\mathsf{nonce})$ to all parties $\triangleright$ Only $\mathsf{HighBits}$; individual $\mathbf{W}_{i}$ stay private
32:
33:Round 3: Response
34:for each party $i\in S$ in parallel do
35:  Compute response: $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ $\triangleright$ No $\lambda_{i}$, no mask
36:  Compute $\mathbf{V}_{i}=\lambda_{i}\cdot c\cdot\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$ $\triangleright$ Masked, for r0-check
37:  Send $(\mathbf{z}_{i},\mathbf{V}_{i})$ to combiner
38:end for
39:
40:Aggregation and r0-Check (see Subsection 3.10)
41:Combiner computes $\mathbf{z}=\sum_{i\in S}\lambda_{i}\cdot\mathbf{z}_{i}\mod q$ $\triangleright$ Lagrange at combiner
42:if $\left\|\mathbf{z}\right\|_{\infty}\geq\gamma_{1}-\beta$ then
43:  return $\bot$ (z-bound abort, retry from Round 1)
@@
45:r0-check: Verify $\left\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\right\|_{\infty}<\gamma_{2}-\beta$
46:if r0-check fails then
47:  return $\bot$ (r0 abort, retry from Round 1)
48:end if
49:Compute $\mathbf{r}=\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d}$ $\triangleright$ $\mathbf{w}-c\mathbf{s}_{2}+c\mathbf{t}_{0}$
50:Compute $c\mathbf{t}_{0}$ using $c\mathbf{s}_{2}$ from r0-check $\triangleright$ See Subsection 3.10
51:Compute hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{r},2\gamma_{2})$
52:if $\mathsf{weight}(\mathbf{h})>\omega$ then
53:  return $\bot$ (hint weight abort)
54:end if
55:return signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
@@
The Round 1 commitments $\mathsf{Com}_{i}$ bind parties to their nonces before seeing others’ values, preventing adaptive attacks. However, we do not publicly verify these commitments during normal signing—the masked values $\mathbf{W}_{i}$ cannot be verified without revealing private mask seeds, and the responses $\mathbf{z}_{i}$ are privacy-protected by the nonce DKG (Theorem 4.4). Instead, we follow an optimistic model: if signing fails (excessive aborts, invalid signatures), the TEE coordinator triggers a blame protocol where parties reveal their commitments for verification. This “commit always, open only on blame” approach minimizes latency on the normal path while enabling accountability. See Section 6.2 for details.
Remark 3.7 (Why $\mathbf{W}_{i}$ Must Not Be Broadcast).
[…]
In Profile P1, the TEE coordinator naturally keeps $\mathbf{W}_{i}$ private. In Profile P3+, the combiner serves as coordinator. In Profile P2 (no designated coordinator), parties broadcast $\mathbf{W}_{i}$ directly and rely on mask hiding (Lemma 4.9, requiring $|S\setminus C|\geq 2$) to protect individual contributions.
@@
We use distinct domain tags ("com", "nonce0", "nonce", "s2", "comm", "pairwise_seed") for all hash function uses to ensure random oracle independence.
@@
When $\gamma_{1}$ is not divisible by $|S|$, rounding slightly reduces the nonce range. This has negligible impact: for ML-DSA-65 with $|S|=17$, the range loss is $2\times 10^{-5}$ per coordinate and does not affect the z-bound check since $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor\geq\gamma_{1}-(|S|-1)>\gamma_{1}-\beta$ (holds for all $|S|\leq\beta+1=197$, covering all practical configurations). See Appendix G for details.
@@ 3.7 Correctness
Proof 3.11.
The aggregated response is $\mathbf{z}=\sum_{i\in S}\lambda_{i}\mathbf{z}_{i}=\sum_{i}\lambda_{i}(\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i})=\sum_{i}\lambda_{i}\mathbf{y}_{i}+c\cdot\sum_{i}\lambda_{i}\mathbf{s}_{1,i}=\mathbf{y}+c\mathbf{s}_{1}$. Both $\{\mathbf{y}_{i}\}$ and $\{\mathbf{s}_{1,i}\}$ are degree-(T1)(T\!-\!1) Shamir shares of $\mathbf{y}$ and $\mathbf{s}_{1}$ respectively, evaluated at the same points, so Lagrange reconstruction yields both correct secrets simultaneously. The masked commitment aggregation gives $\mathbf{W}=\sum_{i}\mathbf{W}_{i}=\sum_{i}(\lambda_{i}\mathbf{w}_{i}+\mathbf{m}_{i}^{(w)})=\mathbf{A}\sum_{i}\lambda_{i}\mathbf{y}_{i}=\mathbf{A}\mathbf{y}=\mathbf{w}$, where mask cancellation follows from Lemma 3.4. Since $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ and $\mathbf{w}=\mathbf{A}\mathbf{y}$, verification proceeds as in single-signer ML-DSA.
@@
(1) Each party computes $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ without applying the Lagrange coefficient $\lambda_{i}$. The combiner applies $\lambda_{i}$ during aggregation: $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}$. While individual $\lambda_{i}\mathbf{z}_{i}$ may have large coefficients, the sum $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ is short (<q/2<q/2), so mod $q$ reduction is correct. (2) Signatures are FIPS 204-compatible: same format, size (3309 bytes for ML-DSA-65), and pass unmodified verifiers. Signer-side differences (Irwin-Hall nonces, nonce DKG, interactive protocol) do not affect security or verification.
@@ 3.8 Nonce Distribution and Security
The aggregated nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ (the sum of each party’s constant-term contribution from the nonce DKG, equivalently $\mathbf{y}=\sum_{i\in S}\lambda_{i}\mathbf{y}_{i}$ by Lagrange reconstruction) follows an Irwin-Hall distribution concentrated around zero. This preserves EUF-CMA security: using smooth Rényi divergence bounds from [Raccoon2024], we compute per-coordinate $R^{\epsilon}_{2,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$. Since nonce coordinates are sampled independently, the full-vector divergence is $R_{2}^{\mathsf{vec}}=(R^{\epsilon}_{2,\mathsf{coord}})^{1280}\approx 2^{5.4}$ for $|S|\leq 17$ ($\approx$3-bit security loss). For $|S|=33$, $R_{2}^{\mathsf{vec}}\approx 2^{75}$ ($\approx$38-bit loss). See Theorem 4.11 and Appendix D.
@@ 4.1 Security Model
Remark 4.1 ($|S|\geq T$ Suffices in Coordinator-Based Profiles).
With Shamir nonce DKG and the coordinator-based communication model (Remark 3.7), Profiles P1 and P3+ require signing sets of size only $|S|\geq T$. The honest party’s response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ has high conditional min-entropy (Theorem 4.4) from the adversary’s view, vastly exceeding the secret key entropy. The key privacy argument rests on what the adversary observes:
  1. 1.
    Response $\mathbf{z}_{h}$: Statistical privacy via nonce DKG (Theorem 4.4). Requires no computational assumptions, holds even with $|S\setminus C|=1$.
  2. 2.
    r0-check shares $\mathbf{V}_{i}$: Never broadcast—processed inside the TEE (P1), as MPC inputs (P2), or via 2PC (P3+). Even with $|S\setminus C|=1$, the adversary does not observe individual $\mathbf{V}_{h}$.
  3. 3.
    Masked commitments $\mathbf{W}_{i}$: Sent only to the coordinator (not broadcast). The coordinator publishes only $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{W},2\gamma_{2})$, which is already implicit in the signature (verifiers reconstruct $\mathbf{w}_{1}$ from $(\mathbf{A},\mathbf{z},c,\mathbf{h})$). This adds zero information to the adversary’s view.
Why individual $\mathbf{W}_{i}$ must stay private.
If $\mathbf{W}_{h}$ were broadcast with $|S\setminus C|=1$, the pairwise mask $\mathbf{m}_{h}^{(w)}$ would be computable by the adversary (all seeds involve corrupted counterparties), exposing $\lambda_{h}\mathbf{A}\mathbf{y}_{h}$. Since $\mathbf{A}\in R_{q}^{k\times\ell}$ with $k>\ell$ is generically injective (under the NTT isomorphism, a random $k\times\ell$ matrix over $\mathbb{Z}_{q}$ has full column rank with probability $\geq 1-\ell/q$), the adversary could recover $\mathbf{y}_{h}$ via linear algebra—this is not a Module-LWE instance because $\mathbf{y}_{h}$ has full-size (not short) coefficients. Combined with $\mathbf{z}_{h}$ (extractable from the public signature), this would yield $\mathbf{s}_{1,h}=c^{-1}(\mathbf{z}_{h}-\mathbf{y}_{h})$, a complete key recovery. The coordinator-only communication model eliminates this attack vector entirely.
Profile-specific analysis.
In P1, the TEE coordinator naturally keeps individual $\mathbf{W}_{i}$ private. In P3+, the combiner serves as coordinator (trusted for privacy under the 1-of-2 CP honest assumption). In P2 (no designated coordinator), parties broadcast $\mathbf{W}_{i}$ directly; mask hiding (Lemma 4.9) protects individual values under the existing $|S\setminus C|\geq 2$ condition, which P2 already requires.
@@ 4.2 Unforgeability
$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\left(O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\tbinom{N}{2}\cdot\epsilon_{\mathsf{PRF}}\right)}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
where $q_{H},q_{s}$ are random oracle and signing queries, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector Rényi divergence with per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$, and $\epsilon_{\mathsf{IH}}<10^{-30}$ (Theorem 4.11). The $\sqrt{R_{2}^{\mathsf{vec}}}$ factor arises from the game hop between Irwin-Hall and uniform nonce distributions via Cauchy-Schwarz; for $|S|\leq 17$, $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$, giving $\sqrt{R_{2}^{\mathsf{vec}}}\approx 2^{2.7}$ ($\approx$3-bit loss). For $|S|=33$, $R_{2}^{\mathsf{vec}}\approx 2^{75}$ ($\approx$38-bit loss). The loss matches the per-coordinate composition inherent to the generic Rényi proof technique [Raccoon2024].
@@
$\mathsf{Game}_{3.5}$ (Uniform Nonces).
Replace the Irwin-Hall nonce distribution with uniform nonces: in each signing session, sample $\mathbf{y}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1},\ldots,\gamma_{1}\}^{n\ell}$ uniform instead of $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ (Irwin-Hall). By the smooth Rényi divergence bound (Theorem 4.11 with $\alpha=2$):
$\Pr[\mathsf{Win}_{3}]\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\Pr[\mathsf{Win}_{3.5}]}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
where $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector Rényi divergence (per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$, $n\ell=1280$), giving $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ for $|S|\leq 17$. The tail probability satisfies $\epsilon_{\mathsf{IH}}<10^{-30}$. This is the only game hop that incurs the Irwin-Hall security loss; individual shares $\mathbf{z}_{j}$ have high min-entropy by Theorem 4.4 (statistical, no computational assumptions).
@@
Starting from uniform nonces (Game 3.5), simulate signing without knowing $\mathbf{s}_{1}$: sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1}+\beta+1,\ldots,\gamma_{1}-\beta-1\}^{n\ell}$, compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$, and program $H(\mu\|\mathbf{w}_{1}):=c$. The masked values $\mathbf{W}_{j}$ are simulated as uniform by Lemma 4.9. By PRF security: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3.5}]|\leq\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
@@
This is an inhomogeneous M-SIS instance (SelfTargetMSIS [KLS18]): the residual $\boldsymbol{\delta}$ arises from rounding in $\mathsf{HighBits}$ and satisfies $\|\boldsymbol{\delta}\|_{\infty}<2\gamma_{2}$, which is absorbed into the M-SIS norm bound. The extracted vector $(\mathbf{v},\Delta c)$ satisfies $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),2)<2\gamma_{1}$, solving the augmented M-SIS instance $\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]$. This is the standard reduction used in the Dilithium/FIPS 204 security proof. Thus: $\Pr[\mathsf{Win}_{4}]\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}$.